OpenStudy (anonymous):

Statistics, constructing upper, lower, and two-sided confidence intervals based on a gamma distribution

OpenStudy (anonymous):

We were explained how to prove part a and b, but if you'd like to give an explanation, i won't hold you back, since it didn't make full sense other than utilizing the method of transformations

OpenStudy (kirbykirby):

One other way you can do a) is to use the CDF method: Let \(Y=2Xi/\beta\) Then \[\large \begin{align} F_Y(y)=P(Y\le y) &=P\left(\frac{2X_i}{\beta}\le y \right)\\ &=P\left(X_i\le\frac{\beta y}{2} \right) \\ &=F_{X_i}\left(\frac{\beta y}{2} \right)\end{align}\] Then differentiate both sides with respect to y: \[\large f_Y(y)=f_{X_i}\left( \frac{\beta y}{2}\right) \cdot \frac{\beta}{2}, \text{(chain rule)}\\ =\large\frac{\beta}{2} \frac{1}{\Gamma(\alpha)\beta^{\alpha}}\left(\frac{\beta y}{2}\right)^{\alpha-1}e^{\frac{-\beta y}{2}/\beta}\\ = \large \frac{1}{\Gamma(\alpha)2^{\alpha}}y^{\alpha-1}e^{-y/2}\] This is the PDF of \(\text{Gamma}(α,2)\) Now just remember that the link between the Gamma and Chi-Square distribution is: \(\large\text{Gamma}\left(\frac{n}{2},2\right)=\chi^2(n)\), so I hope you see that you get a \(\chi^2(2\alpha)\) distribution (note you can write your gamma distribution as \(\text{Gam}(2\alpha/2,2)\) ). For part b), I hope you remember that the sum of independent chi-squared distributions is a chi-square distribution with all the individual degrees of freedom added. You can easily show this via moment generating functions

OpenStudy (kirbykirby):

recall that the MGF for \(X\sim\text{Gamma}(\alpha,\beta)\) is \[M_X(t)=\frac{1}{(1-\beta t)^{\alpha}}\] So for \(\chi^2(2\alpha)=\text{Gam}(\alpha,2)\): \[M_X(t)=\frac{1}{(1-2t)^{\alpha}}\] Now, using the same variable Y as defined above, and using the random variable V defined in the question, Recall for independent random variables, \(V=Y_1+Y_2+\ldots +Y_n \), \[\large\begin{align} M_V(t)&=\prod_{i=1}^nM_{Y_i}(t)\\ &=\prod_{i=1}^n\frac{1}{(1-2t)^{\alpha}}\\ &=\frac{1}{(1-2t)^{n\alpha}}\end{align}\] This is the MGF of \(\text{Gam}(n\alpha,2)=\chi^2(2n\alpha)\), so \(V\) has this distribution (I am not sure why the question argue though that you have \(n\) degrees of freedom? when I believe it should be \(2n\alpha\) degrees of freedom, because the parameter of the Chi-squre distribution are known as degrees of freedom).

OpenStudy (kirbykirby):

We can use V as a pivot for \(\beta\) because the distribution of V does not have the parameter \(\beta\) in it. This is from the definition of a pivot, which is a function whose probability distribution does not depend on unknown parameters (even though \(\alpha\) is unknown, it is a pivot for \(\beta\) since \(\beta\) is not in the distribution of V).

OpenStudy (kirbykirby):

For the confidence interval.. first observe that \(\sum_{i=1}^nX_i =n\bar{X}\) So with \(n=5\) and \(\bar{X}=5.39\), \(\sum_{i=1}^5\bar{X}=26.95\) Also, with \(\alpha=2\), \(\chi^2(2n\alpha)=\chi^2(20)\) So to construct your (exact) confidence interval, you need to use the pivotal quantity (pivot) since it doesn't depend on unknown parameters. A general way of writing a \(100p\%\) 2-sided confidence interval with a pivotal quantity \(Q(X;\theta)\), where \(\theta\) is just the unknown parameter(s) of your distribution of \(X\), is first writing \[P( q_1\le Q(X;\theta)\le q_2)=p\], then the confidence interval is just \( [q_1, q_2]\), where \(q_1\) and \(q_2\) are the lower and upper quantiles of your pivotal quantity. For a 95% 2-sided confidence interval in your case would be first writing: \[\large \begin{align}P\left(\chi^2_{0.025}(20)\le\frac{2}{\beta} \sum_{i=1}^5X_i \le\chi^2_{0.975}(20)\right)&=0.95 \\ P\left( \chi^2_{0.025}(20)\le \frac{2}{\beta}(26.95)\le \chi^2_{0.975}(20)\right)&=0.95 \\ P\left(\chi^2_{0.025}(20)\le \frac{53.9}{\beta}\le \chi^2_{0.975}(20)\right)&=0.95 \\ P(\left( \frac{1}{\chi^2_{0.025(20)}}\ge\frac{\beta}{53.9}\ge\frac{1}{\chi^2_{0.975(20)}}\right)&=0.95\\ P\left( \frac{53.9}{\chi^2_{0.025(20)}}\ge\beta\ge\frac{53.9}{\chi^2_{0.975(20)}}\right)&=0.95\\ P\left(\frac{53.9}{\chi^2_{0.975(20)}}\le \beta\le \frac{53.9}{\chi^2_{0.025(20)}} \right)&=0.95\end{align}\] So a 95% 2-sided CI for \(\beta\)is: \[\large \left[\frac{53.9}{\chi^2_{0.975(20)}}, \frac{53.9}{\chi^2_{0.025(20)}} \right] \] those values are just the quantiles which you can find in a chi-square table with 20 degrees of freedom (just be careful whether the table is showing you CDF values, or tail values). Now a one-sided confidence interval is similar as above, but only one inequality is needed and you would need to use quantiles like \(\large\chi^2_{0.05}(20)\) and \(\large\chi^2_{0.95}(20)\). Although I have rarely used one-sided confidence intervals, I believe an upper 95% confidence interval would be based on the result \(\large P(Q(X;\theta)\ge \chi^2_{0.05}(20))=0.95\) and the CI should be in the form \([q_1,\infty)\) a lower 95% confidence interval would be based on the result \(\large P\left(Q(X;\theta)\le \chi^2_{0.95}(20)\right)= 0.95\) and the CI should be in the form \([0,q_2]\) since the chi-square distribution has support \(0\le x < \infty\) If you draw all three confidence intervals, the interval in which all 3 overlap should be \([q_1,q_2]\) when you find those values. ----- Also, I just want to add, if you're wondering why doing all the steps in a) and b) were necessary... to find a confidence interval for \(\beta\) based on the gamma distribution given can't be done (at least for finding EXACT confidence intervals), because the gamma distribution given is dependent on \(\beta\) and is thus not a pivotal quantity. We eventually found a new gamma distribution which was a pivotal quantity because its distribution was not dependent on \(\beta\). And then we converted the gamma into a Chi-square distribution because we can use chi-square tables that are readily available to find the numeric values of the quantiles as we did above.

OpenStudy (anonymous):

Thank you so much, i apologize for not commenting or anything when you were answering the first parts, i was in my lectures and did not access the internet during them, and after looking at everything i wanted to see if i really understood so i went and completed another similar problem without issues. So again, thank you so much :D your explanations never fail to help me understand

OpenStudy (kirbykirby):

Your welcome, and awesome :)!